A canoe on a river flowing 0.50 𝑚/𝑠 east has a velocity of 0.40 𝑚/𝑠 southeast relative to the earth. Find the magnitude and the direction of the canoe’s velocity relative to the river. 


1
Expert's answer
2021-09-25T13:09:27-0400

Given:

"{\bf u}=0.50\hat i\:\rm m/s"

"{\bf V}=\frac{0.40}{\sqrt{2}}(\hat i-\hat j)\:\rm m/s"


The velocity of a canoe relative to the earth

"{\bf V}={\bf u}+{\bf v}"

Hence, the velocity of a canoe relative to the river

"{\bf v}={\bf V}-{\bf u}"

"=\frac{0.40}{\sqrt{2}}(\hat i-\hat j)-0.50\hat i=-0.217\hat i-0.283\hat j"

Magnitude:

"v=\sqrt{0.217^2+0.283^2}=0.36\:\rm m/s"

Direction:

"\theta=\tan^{-1}\frac{-0.283}{-0.217}=53^{\circ}\quad \rm S\;of\; W"